One's intelligence quotient or IQ varies directly as a person's mental age and inversely as that person's chronological age. A person with a mental age of 15 and a chronological age of 25 has an IQ of 120. What is the chronological age of a person with a mental age of 20 and an IQ of 100? The chronological age is what?
Answer #1

Given that,

IQ varies directly as a person's mental age

I.Q\alpha MentalAge

IQ varies inversely as that person's chronological age.

I.Q\alpha \frac{1}{ChronologicalAge}

Combining both

I.Q\alpha \frac{MentalAge}{ChronologicalAge}

Equating to a constant k

I.Q=k \frac{MentalAge}{ChronologicalAge}

Given that a person with a mental age of 15 and a chronological age of 25 has an IQ of 120

120=k \frac{15}{25}

k=120\times \frac{25}{15}

Given another mental age of 20 and an IQ of 100

I.Q=k \frac{MentalAge}{ChronologicalAge}

100=k\frac{20}{ChronologicalAge}

Put k=120\times \frac{25}{15}

100=120\times \frac{25}{15}\times\frac{20}{ChronologicalAge}

100=120\times \frac{5}{3}\times\frac{20}{ChronologicalAge}

100=40\times \frac{5}{1}\times\frac{20}{ChronologicalAge}

ChronologicalAge=40\times \frac{5}{1}\times\frac{20}{100}

ChronologicalAge=4\times \frac{5}{1}\times\frac{2}{1}

ChronologicalAge=40

Therefore,

Chronological age of a person with a mental age of 20 and an IQ of 100 is 40